An object moving at speed of 20ms-1 loses its speed by 3ms-2 . How does it travel before it stops?
1
Expert's answer
2019-07-16T09:02:51-0400

We can find the distance traveled by the object before it stops from the kinematic equation:


"v_f^2 = v_i^2 + 2ad,"

here, "v_i = 20 m\/s" is the initial velocity of the object, "v_f = 0" is the final velocity of the object, "a = -3 m\/s^2" is the deceleration of the object and "d" is the distance traveled by the object before it stops.

Then, from this formula we can calculate the distance traveled by the object before it stops:


"d = \dfrac{-v_i^2}{2a} = \dfrac{-(20 \dfrac{m}{s})^2}{2 \cdot (-3 \dfrac{m}{s^2})} = 67 \dfrac{m}{s}."

Answer:

"d = 67 \dfrac{m}{s}."